# Ding Chen1  Qingchen Yu211footnotemark: 1  Pengyuan Wang211footnotemark: 1  Wentao Zhang322footnotemark: 2  &Bo Tang2  Feiyu Xiong2  Xinchi Li1  Minchuan Yang1  Zhiyu Li [^2] footnotemark: 2

### v1

@@ C.2 Original Model Details
Appendix D Prompts
@@
D.1 Prompts for Generating LLM Responses
The prompt templates used to generate LLM responses are illustrated in Figures 4 to 7. Each template consists of four fields that need to be populated: "task_type", "task_description", "examples", and "question". The "task_type" and "task_description" fields are determined based on the type of question. For instance, for questions from the GPQA dataset, "task_type" is set to "multidisciplinary question", and "task_description" is set to "Please choose the answer from options A to D, corresponding to the question." During dataset preprocessing, we design appropriate "task_type" and "task_description" values for each dataset. The "examples" field is filled according to the selected prompting strategy, either 0-shot or 5-shot. In the 0-shot setting, this field is left empty, while in the 5-shot setting, it is populated with five example question-answer pairs that are similar to the target "question". The "question" field contains the specific query to be answered by the LLM. Examples of the "examples" and "question" fields are shown in Figures 8 and 9, respectively.
@@ D.1 Prompts for Generating LLM Responses
***** Start In-Context Examples ***** Q: A late game rally by Washington led them to the Eagles’ 26 yard line. A shot to the end zone by Robert Griffin III would be intercepted by Brandon Boykin, clinching an Eagles win. The Eagles would move to 6-5. This is the Eagles first win at Lincoln Financial Field since Week 4 of the 2012 season, because prior to this game, the Eagles had never won a game in their home stadium in 414 days since that same week, snapping a 10-game losing streak at home with this win. How many more wins than losses did the Eagles have after this game? A: The answer is 1. Q: The population of Sevastopol proper is 418,987 (01.01.16), making it the largest in the Crimean Peninsula. The citys agglomeration has about 600,000 people (2015). According to the Ukrainian Census (2001), the ethnic groups of Sevastopol include Russians (71.6%), Ukrainians (22.4%), Belarusians (1.6%), Tatars (0.7%), Crimean Tatars (0.5%), Armenians (0.3%), Jews (0.3%), Moldovans (0.2%), and Azerbaijani people (0.2%). Which ethnic has a higher percentage of the population in Sevastopol: Russians or Armenians? A: The answer is Russians. Q: the most common crimes in the ACT are property related crimes, unlawful entry with intent and motor vehicle theft. They affected 2,304 and 966 people (580 and 243 per 100,000 persons respectively). Homicide and related offences—murder, attempted murder and manslaughter, but excluding driving causing death and conspiracy to murder—affect 1.0 per 100,000 persons, which is below the national average of 1.9 per 100,000. Rates of sexual assault (64.4 per 100,000 persons) are also below the national average (98.5 per 100,000). Which was there a higher national average for, homicide and related offences or sexual assault? A: The answer is sexual assault. Q: In the county, the population was spread out with 21.7% under the age of 18, 8.5% from 18 to 24, 26.9% from 25 to 44, 27.7% from 45 to 64, and 15.0% who were 65 years of age or older. The median age was 40 years. For every 100 females, there were 94.4 males. For every 100 females age 18 and over, there were 98.7 males. How many percent were not from 45 to 64? A: The answer is 72.3. Q: The median age in the city was 35.1 years. 24.2% of residents were under the age of 18; 7.9% were between the ages of 18 and 24; 33.8% were from 25 to 44; 24.6% were from 45 to 64; and 9.5% were 65 years of age or older. The gender makeup of the city was 48.6% male and 51.4% females. How many more people, in terms of percentage, were in the largest age group compared to the second smallest? A: The answer is 24.3. ***** End In-Context Examples *****
@@
Q: Let $ABCD$ be a tetrahedron such that $AB=CD=\sqrt{41}$, $AC=BD=\sqrt{80}$, and $BC=AD=\sqrt{89}$. There exists a point $I$ inside the tetrahedron such that the distances from $I$ to each of the faces of the tetrahedron are all equal. This distance can be written in the form $\frac{m\sqrt{n}}{p}$, where $m$, $n$, and $p$ are positive integers, $m$ and $p$ are relatively prime, and $n$ is not divisible by the square of any prime. Find $m+n+p$. A:
@@
D.2 Prompts for GPT-4o Annotation
The prompt templates used for annotating the collected LLM question-answer pairs with GPT-4o during the construction of the VAR dataset are shown in Figures 10 and 11. Both of these prompt templates employ the Chain-of-Thought (CoT) strategy to ensure the accuracy of the annotations generated by GPT-4o.
@@ D.2 Prompts for GPT-4o Annotation
You are a diligent and precise assistant tasked with evaluating the correctness of responses. Think step by step as you make your evaluation. We request your feedback on whether the model’s response correctly answers the user question above. Follow these steps to make your evaluation: 1. Think step by step: Read the user question carefully. 2. Think step by step: Review the reference answer and understand the key points it covers. 3. Think step by step: Compare the model’s answer with the reference answer. 4. Think step by step: Determine if the model’s answer addresses the key points in the reference answer and correctly answers the question. - First, provide your reasoning in detail. Then, clearly state your judgment as either "Correct" or "Incorrect." Please present your response in the following JSON format: { "reasoning": "Your step-by-step reasoning here.",
@@
D.3 Prompts for Data Augmentation
In constructing the VAR dataset, two prompt templates used to guide GPT-4o in augmenting mathematical question samples are presented in Figures 12 and 13.
You are an expert in mathematical calculations and data expressions. You are
required to provide different equivalent forms of the standard answer for the
following math problem.
@@ D.3 Prompts for Data Augmentation
Problem: A volcano erupts and spews ash into the sky. The ash cloud spreads out
in a diameter eighteen times as far as the distance it shot up into the sky. If
the ashes erupted three hundred feet into the sky, what was the radius of the
ash cloud in feet?
@@
    "answer3": "2.7 $\times$ 10^3",
    "answer4": "\$2.7 $\times$ 10^3\$",
@@
2. Each different form must be equivalent to the standard answer, i.e., it
should still be a correct and valid answer.
3. You may use LaTeX, scientific notation, or other standard mathematical
expressions.
@@ 4. Please follow the JSON format below for the output:
You are an expert in mathematical calculations and data expressions. For an
answer to a specific mathematical problem, you are required to provide
equivalent and different expressions of the mathematical result.
@@
    "answer1": "The answer is 2.7×10^3 feet.",
@@
    "answer3": "The answer is 2.7 $\times$ 10^3 feet.",
[…]
@@
1. You need to provide 3 to 5 different expressions, each replacing the
mathematical result with an equivalent and different form.
2. Each expression must be exactly equivalent to the target answer to ensure
its correctness.
3. You can use LaTeX, scientific notation, or other standard mathematical
formats.
@@ 4. Please output the result in the following JSON format:
D.4 Prompts for Judge Model
@@ D.5 Prompts for xVerify
Appendix E Supplementary Experimental Results
E.1 Evaluation Accuracy Results of All xVerify Models
Tables 19 and 20 present the performance of all 14 xVerify models on the test set and the generalization set, respectively. Overall, each xVerify model achieves an F1 score and accuracy exceeding 96.5% on the test set and over 95.52% on the generalization set. These results demonstrate not only the effectiveness of the xVerify models for evaluation tasks but also the high quality of the VAR dataset.
A comparison between the results on the two datasets shows that the performance on the generalization set experiences a slight decline relative to the test set, with the decrease not exceeding 1.6%. Moreover, models with larger parameter sizes exhibit smaller performance drops. This indicates that the xVerify models possess strong generalization capabilities, which further improve with an increase in parameter scale. Additionally, it is observed across both datasets that while the performance of xVerify models generally enhances with the increment of parameter size, beyond a certain threshold, further increases in parameter scale do not lead to additional performance gains.
Table 19: Evaluation Accuracy Results on the Test Set: All xVerify Models. The best performance in each column is shown in bold, and the second-best performance is underlined.
xVerify Model Multiple Choice Math Short Answer Classification Total
@@ E.1 Evaluation Accuracy Results of All xVerify Models
Table 20: Evaluation Accuracy Results on the Generalization Set: All xVerify Models. The best performance in each column is shown in bold, and the second-best performance is underlined.
Multiple Choice Math Short Answer Classification Total
xVerify Model F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc.
@@
E.2 Computational Efficiency and Operational Cost of xVerify and Judge Models

### v2

@@ C.2 Original Model Details
Appendix D RL Training Details
This section will further present additional information about RL training.
D.1 RL Training Hyperparameters
We train our models using the veRL (sheng2024hybridflow) Framework. The hyperparameters for implementing RL are presented in Table 20. When implementing baseline methods, we use the same hyperparameters. The training is carried out on multiple GPU servers.
Hyperparameter Setting
Epoch 15
Learning Rate 1.0e-6
Train Batch Size 1024
Maximum Rollout Length 8192
KL Loss Coefficient 1e-3
Entropy Coefficient 0.0
Temperature 1.0
Rollouts per Prompt 5
Table 20: Hyperparameter settings for RL training.
D.2 Details of RL Training and Generalization Sets
When training policy, we use xVerify as the reward model. To ensure that the evaluation truly reflects the generalization capability of xVerify, we select all training data from the generalization set, thereby avoiding data contamination. The training set comprises 20,400 samples. The specific composition is shown in Table 21. The generalization set comprises 1,750 samples as shown in Table 22.
Dataset Sample Counts
MMLU-Pro 3400
MATH 3400
DROP 3400
GSM8K 3400
AgNews 3400
ARC 3400
Table 21: Number of samples from each dataset in RL training set.
Dataset Sample Counts
MATH-500 470
MMLU-Pro 250
DROP 250
GPQA 250
Amazon 250
CMMLU 250
CHID 250
Table 22: Number of samples from each dataset in RL generalization set.
Appendix E Prompts
@@
E.1 Prompts for Generating LLM Responses
The prompt templates used to generate LLM responses are illustrated in Figures 5 to 8. Each template consists of four fields that need to be populated: "task_type", "task_description", "examples", and "question". The "task_type" and "task_description" fields are determined based on the type of question. For instance, for questions from the GPQA dataset, "task_type" is set to "multidisciplinary question", and "task_description" is set to "Please choose the answer from options A to D, corresponding to the question." During dataset preprocessing, we design appropriate "task_type" and "task_description" values for each dataset. The "examples" field is filled according to the selected prompting strategy, either 0-shot or 5-shot. In the 0-shot setting, this field is left empty, while in the 5-shot setting, it is populated with five example question-answer pairs that are similar to the target "question". The "question" field contains the specific query to be answered by the LLM. Examples of the "examples" and "question" fields are shown in Figures 9 and 10, respectively.
@@ E.1 Prompts for Generating LLM Responses
***** Start In-Context Examples ***** Q: A late game rally by Washington led them to the Eagles’ 26 yard line. A shot to the end zone by Robert Griffin III would be intercepted by Brandon Boykin, clinching an Eagles win. The Eagles would move to 6-5. This is the Eagles first win at Lincoln Financial Field since Week 4 of the 2012 season, because prior to this game, the Eagles had never won a game in their home stadium in 414 days since that same week, snapping a 10-game losing streak at home with this win. How many more wins than losses did the Eagles have after this game? A: The answer is 1. Q: The population of Sevastopol proper is 418,987 (01.01.16), making it the largest in the Crimean Peninsula. The citys agglomeration has about 600,000 people (2015). According to the Ukrainian Census (2001), the ethnic groups of Sevastopol include Russians (71.6%), Ukrainians (22.4%), Belarusians (1.6%), Tatars (0.7%), Crimean Tatars (0.5%), Armenians (0.3%), Jews (0.3%), Moldovans (0.2%), and Azerbaijani people (0.2%). Which ethnic has a higher percentage of the population in Sevastopol: Russians or Armenians? A: The answer is Russians. Q: the most common crimes in the ACT are property related crimes, unlawful entry with intent and motor vehicle theft. They affected 2,304 and 966 people (580 and 243 per 100,000 persons respectively). Homicide and related offences—murder, attempted murder and manslaughter, but excluding driving causing death and conspiracy to murder—affect 1.0 per 100,000 persons, which is below the national average of 1.9 per 100,000. Rates of sexual assault (64.4 per 100,000 persons) are also below the national average (98.5 per 100,000). Which was there a higher national average for, homicide and related offences or sexual assault? A: The answer is sexual assault. Q: In the county, the population was spread out with 21.7% under the age of 18, 8.5% from 18 to 24, 26.9% from 25 to $4$, 27.7% from 45 to 64, and 15.0% who were 65 years of age or older. The median age was 40 years. For every 100 females, there were 94.4 males. For every 100 females age 18 and over, there were 98.7 males. How many percent were not from 45 to 64? A: The answer is 72.3. Q: The median age in the city was 35.1 years. 24.2% of residents were under the age of 18; 7.9% were between the ages of 18 and 24; 33.8% were from 25 to 44; 24.6% were from 45 to 64; and 9.5% were 65 years of age or older. The gender makeup of the city was 48.6% male and 51.4% females. How many more people, in terms of percentage, were in the largest age group compared to the second smallest? A: The answer is 24.3. ***** End In-Context Examples *****
@@
Q: Let ABCDABCD be a tetrahedron such that $AB=CD=\sqrt{41}$, $AC=BD=\sqrt{80}$, and $BC=AD=\sqrt{89}$. There exists a point II inside the tetrahedron such that the distances from II to each of the faces of the tetrahedron are all equal. This distance can be written in the form $\frac{m\sqrt{n}}{p}$, where $m$, $n$, and $p$ are positive integers, $m$ and $p$ are relatively prime, and $n$ is not divisible by the square of any prime. Find $m+n+p$. A:
@@
E.2 Prompts for GPT-4o Annotation
The prompt templates used for annotating the collected LLM question-answer pairs with GPT-4o during the construction of the VAR dataset are shown in Figures 11 and 12. Both of these prompt templates employ the Chain-of-Thought (CoT) strategy to ensure the accuracy of the annotations generated by GPT-4o.
You are a diligent and precise assistant tasked with evaluating the correctness of responses. Think step by step as you make your evaluation.
You will receive a question, an output sentence, and the correct answer. Your task is to determine if the output sentence accurately answers the question based on the provided correct answer. Think step by step and respond with either [Correct] or [Incorrect]. - Special considerations: 1. **Multiple Answers**: If the output contains multiple answers, evaluate whether later answers modify or correct earlier ones. In such cases, compare the final answer with the correct answer. If the final answer is unclear or incorrect, respond with [Incorrect]. 2. **Mathematical Problems**: If the formats differ but the answers are mathematically equivalent, respond with [Correct]. 3. **Explicit Options**: If the question provides explicit candidate answers, the output will be considered correct if it clearly indicates the correct option’s code or the correct option’s content. 4. **No Explicit Options**: If the question does not provide explicit options, the output must align with the correct answer in content and meaning to be considered [Correct]. Please present your response in the following JSON format: { "reasoning": "Your step-by-step reasoning here.",
@@ E.2 Prompts for GPT-4o Annotation
You are a diligent and precise assistant tasked with evaluating the correctness of responses. Think step by step as you make your evaluation.
We request your feedback on whether the model’s response correctly answers the user question above. Follow these steps to make your evaluation: 1. Think step by step: Read the user question carefully. 2. Think step by step: Review the reference answer and understand the key points it covers. 3. Think step by step: Compare the model’s answer with the reference answer. 4. Think step by step: Determine if the model’s answer addresses the key points in the reference answer and correctly answers the question. - First, provide your reasoning in detail. Then, clearly state your judgment as either "Correct" or "Incorrect." Please present your response in the following JSON format: { "reasoning": "Your step-by-step reasoning here.",
@@
E.3 Prompts for Data Augmentation
In constructing the VAR dataset, two prompt templates used to guide GPT-4o in augmenting mathematical question samples are presented in Figures 13 and 14.
You are an expert in mathematical calculations and data expressions.
You are required to provide different equivalent forms of the standard
answer for the following math problem.
@@ E.3 Prompts for Data Augmentation
Problem: A volcano erupts and spews ash into the sky. The ash cloud
spreads out in a diameter eighteen times as far as the distance it
shot up into the sky. If the ashes erupted three hundred feet into the
sky, what was the radius of the ash cloud in feet?
@@
    "answer3": "2.7 $\times$ 10ˆ3",
    "answer4": "\$2.7 $\times$ 10ˆ3\$",
@@
2. Each different form must be equivalent to the standard answer, i.e.,
it should still be a correct and valid answer.
3. You may use LaTeX, scientific notation, or other standard
mathematical expressions.
@@ 4. Please follow the JSON format below for the output:
You are an expert in mathematical calculations and data expressions.
For an answer to a specific mathematical problem, you are required to
provide equivalent and different expressions of the mathematical
result.
@@
    "answer1": "The answer is 2.7×10ˆ3 feet.",
@@
    "answer3": "The answer is 2.7 $\times$ 10ˆ3 feet.",
    "answer4": "The answer is \$2.7 $\times$ 10ˆ{3}\$ feet.",
@@
1. You need to provide 3 to 5 different expressions, each replacing
the mathematical result with an equivalent and different form.
2. Each expression must be exactly equivalent to the target answer to
ensure its correctness.
3. You can use LaTeX, scientific notation, or other standard
mathematical formats.
@@ 4. Please output the result in the following JSON format:
E.4 Prompts for Judge Model
@@ E.5 Prompts for xVerify
Appendix F Supplementary Experimental Results
@@
Table 23: Evaluation Accuracy Results on the Test Set: All xVerify Models and Judge Models. The best performance in each column is shown in bold, and the second-best performance is underlined.
F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc.
@@
Table 24: Evaluation Accuracy Results on the Generalization Set: All xVerify Models and Judge Models. The best performance in each column is shown in bold, and the second-best performance is underlined.
F.1 Evaluation Accuracy Results of All xVerify Models and Judge Models
Tables 23 and 24 present the performance of all xVerify models and judge models on the test set and generalization set, respectively. Overall, each xVerify model achieves an F1 score and accuracy exceeding 96.5% on the test set, and an accuracy greater than 95.52% on the generalization set. These results not only demonstrate the effectiveness of xVerify models in the evaluation task—consistently outperforming all other judge models—but also underscore the high quality of the VAR dataset.
A comparison of the results across the two datasets reveals that the performance of xVerify models on the generalization set exhibits a slight decline relative to the test set, with a maximum drop of no more than 1.6%. Moreover, xVerify models with larger parameter sizes tend to show a smaller performance degradation, indicating strong generalization capabilities that further improve with model scale. Additionally, it is observed across both datasets that while the performance of xVerify models generally enhances with the increment of parameter size, beyond a certain threshold, further increases in parameter scale do not lead to additional performance gains.
@@ F.1 Evaluation Accuracy Results of All xVerify Models and Judge Models
F.2 Computational Efficiency and Operational Cost of xVerify and Judge Models
@@
F.3 The Consistency between xVerify and Human Evaluation
To evaluate the effectiveness of the xVerify, we conduct a comparative analysis of its consistency against human-annotated results. We evaluate the consistency between different models and human judgments across multiple evaluation datasets. The evaluation datasets and models utilized in this experiment are consistent with those employed in RL experiments (details in Appendix D). Figure 21 illustrates that xVerify consistently achieves a high degree of alignment with human evaluation across multiple datasets and in diverse scenarios.
Figure 21: The Consistency between xVerify and Human Evaluation
F.4 Learning Curves of Qwen2.5-7B and Llama3.1-8B in Reinforcement Learning
Figures 22 and 23 present the learning curves of Qwen2.5-7B and Llama3.1-8B during reinforcement learning training. It is evident from these curves that using xVerify as the reward model results in higher initial accuracy, indicating that xVerify provides more precise reward signals at the beginning of training. Moreover, the learning curves with xVerify converge faster, demonstrating that it delivers more accurate and reliable feedback, thereby significantly improving training efficiency. This is particularly important for larger-scale and broader reinforcement learning fine-tuning, where both training efficiency and reward signal quality are critical.
Figure 22: Learning Curves of Qwen2.5-7B in Reinforcement Learning
Figure 23: Learning Curves of Llama3.1-8B in Reinforcement Learning
F.5 Full Fine-Tuning vs. QLoRA for xVerify Models
Model Fine-tuning Type Test Set Generalization Set
F1 Acc. F1 Acc.
Qwen2.5-0.5B-Instruct QLoRA 96.69% 96.85% 95.29% 95.53%
Full 95.90% 96.10% 92.64% 93.03%
LLaMA-3.2-3B-Instruct QLoRA 97.02% 97.14% 96.11% 96.27%
Full 95.44% 95.67% 93.10% 93.41%
Table 27: Performance comparison of QLoRA and full fine-tuning on xVerify models.
To quickly explore the impact of different fine-tuning strategies on xVerify models, we conducted a small-scale experiment comparing full fine-tuning with QLoRA on two representative models: Qwen2.5-0.5B-Instruct and LLaMA-3.2-3B-Instruct. Both models were fine-tuned on the VAR dataset, and their performance was evaluated on the test set as well as a held-out generalization set.
Table 27 summarizes the results. As shown, full fine-tuning achieves slightly lower overall performance compared to QLoRA. The decline is more pronounced on the generalization set, suggesting that full fine-tuning may be more prone to overfitting. In contrast, QLoRA maintains a better balance between test performance and generalization capability.
These findings indicate that, although full fine-tuning could in principle offer marginal gains under some conditions, in this small-scale exploration it does not outperform QLoRA. Additionally, QLoRA provides substantial computational savings, making it a more efficient choice for training instruction-tuned models. Overall, these results highlight that QLoRA is a practical and effective fine-tuning strategy for xVerify.
F.6 Comparison of xVerify models with their base models
Model Test Set Generalization Set
F1 Acc. F1 Acc.
Qwen2.5-0.5B-Instruct 70.55% 66.71% 72.38% 68.78%
xVerify-0.5B-I 96.69% 96.85% 95.29% 95.53%
LLaMA-3.2-1B-Instruct 37.15% 53.95% 41.87% 56.09%
xVerify-1B-I 96.77% 96.91% 95.43% 95.62%
LLaMA-3.2-3B-Instruct 85.22% 84.99% 83.79% 83.32%
xVerify-3B-Ia 97.02% 97.14% 96.11% 96.27%
Gemma-2-9B-it 83.31% 81.89% 82.20% 80.61%
xVerify-9B-I 97.19% 97.29% 96.23% 96.38%
Table 28: Comparison of xVerify models with their base models on the test and generalization sets.
To assess the effectiveness of VAR-based fine-tuning, we conducted a systematic evaluation of four xVerify variants and their corresponding base models on both the test and generalization sets, using identical evaluation settings. As shown in Table 28, the fine-tuned xVerify models achieve over 10 percentage points of gain in both F1 and accuracy compared to their base models. Notably, xVerify-0.5B-I and xVerify-1B-I each deliver improvements of 20%–50% on these metrics. These results demonstrate that VAR-based fine-tuning substantially enhances the evaluation capabilities of the base models.
F.7 Evaluation without the original question
Dataset Model Setting F1 Acc.
Test xVerify-0.5B-I with question 96.69% 96.85%
w/o question 96.53% 96.73%
xVerify-3B-Ia with question 97.02% 97.14%
w/o question 96.09% 96.30%
Generalization xVerify-0.5B-I with question 95.29% 95.53%
w/o question 95.05% 95.33%
xVerify-3B-Ia with question 96.11% 96.27%
w/o question 95.55% 95.79%
Table 29: Performance of xVerify models with and without access to the original question.
In real-world evaluation scenarios, it is often the case that only the LLM’s response and the reference answer are available, without access to the original question. This setting arises, for example, when xVerify is used as a reward model in reinforcement learning.
To assess whether xVerify can maintain strong accuracy and generalization under these conditions, we evaluated xVerify-0.5B-I and xVerify-3B-Ia on both the test set and the generalization set, with and without providing the original questions. In the modified setup, we removed only the parts of the prompt referencing the question, while keeping all other configurations identical.
As shown in Table 29, both models exhibit virtually no performance drop across datasets, demonstrating that xVerify retains robust decision-making capability and generalization even in the absence of the original question. This suggests that xVerify’s core judgment mechanism relies primarily on the equivalence between the model’s output and the reference answer, while the original question serves only as auxiliary context rather than essential input.